# Polaritonic Control of Blackbody Infrared Radiative Dissociation

### v2

@@ IV   Conclusions
We reported microcavity-assisted dissociation rate enhancements of $O(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system. Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced diatomic thermal radiative dissociation changes via the most energetic bound state.
Our work focused on how optical microcavities can modify BIRD reactions of diatomic molecules via a particular mechanism. Despite focusing on these specific systems, the main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly. We expect the analysis introduced here will be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.

### v3

@@ IV   Conclusions
We reported microcavity-assisted dissociation rate enhancements of $O(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 25~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few tens of microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a Rabi frequency dependent enhanced photon-weighted polariton DOS at a targeted frequency, which enables the increase of transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The largest enhancements are obtained when photon modes highly off resonant with the host molecular system are assumed to form polaritons, yet as we showed, sizable polariton-assisted BIRD rate enhancements remain achievable when these highly off resonant interactions are cutoff to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors here reported must be viewed as upper bounds for selected polariton-induced diatomic thermal radiative dissociation via the most energetic bound state.
The inclusion of bound-to-continuum transitions in our analysis alters the dissociation dynamics compared to the case where the reaction proceeds via a doorway state (the highest energy bound state). Dissociation via radiative bound-continuum transitions reveals a broader and smoother enhancement profile under microcavity length variation in the weak coupling regime. While the magnitude of the enhancement remains similar, the key difference lies in the involvement of a much greater number of vibrational transitions relative to the doorway mechanism. In the polariton-assisted case, relative rate enhancement is generally smaller, but it no longer requires precisely matching a particular diatomic transition frequency to the host (strongly coupled) material. Instead, the multiplicity of direct dissociation pathways makes polariton-enhanced BIRD inherently more robust, enabling broadband enhancement across a wide spectrum of host material frequencies $\omega_{M}$.
Our work focused on how optical microcavities can modify diatomic blackbody infrared radiative dissociation. While our analysis ignores collisional energy transfer, and therefore may be viewed as providing upper bounds to the microcavity effect on diatomic BIRD rates, the main qualitative conclusions expressed above are expected to apply broadly to systems where thermal radiation acts as a relevant energy source. We anticipate that the framework developed here will aid in understanding and controlling thermal radiative-induced dissociation processes in more complex, polyatomic systems.
Supporting Information. Results for the HF and LiH molecules, temperature analysis, and additional computational details, including derivations, model parameters, and sensitivity analysis.